OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (7, 0) and a directrix at x = -7.

OpenStudy (anonymous):

@aum please help

OpenStudy (anonymous):

@dumbcow @Hero

OpenStudy (anonymous):

a min...

OpenStudy (anonymous):

y= a(x-h)2+k plug in values...

OpenStudy (anonymous):



OpenStudy (aum):

Let (x,y) be any point on the parabola. The property of a parabola is that any point on the parabola is equidistant from the focus and the directrix. It means, distance of point (x,y) from (7,0) = distance of point (x,y) from the directrix which here is a vertical line x = -7. You can square both sides and equate the square of the distances. Square of the distance (x,y) from (7,0) = (x-7)^2 + (y-0)^2 Square of the distance (x,y) from x = -7 is: (x + 7)^2 Equate the two: (x-7)^2 + y^2 = (x+7)^2 y^2 = (x+7)^2 - (x-7)^2 y^2 = (x+7+x-7)(x+7-x+7) y^2 = (2x)(14) y^2 = 28x

OpenStudy (anonymous):

i thought it was x=1/28y^2

OpenStudy (anonymous):

because y^2=28x isnt one of my answer choices